### v1

# IDOL: Instant Photorealistic 3D Human Creation from a Single Image

Abstract
Creating a high-fidelity, animatable 3D full-body avatar from a single image is a challenging task due to the diverse appearance and poses of humans and the limited availability of high-quality training data. To achieve fast and high-quality human reconstruction, this work rethinks the task from the perspectives of dataset, model, and representation. First, we introduce a large-scale HUman-centric GEnerated dataset, HuGe100K, consisting of 100K diverse, photorealistic sets of human images. Each set contains 24-view frames in specific human poses, generated using a pose-controllable image-to-multi-view model. Next, leveraging the diversity in views, poses, and appearances within HuGe100K, we develop a scalable feed-forward transformer model to predict a 3D human Gaussian representation in a uniform space from a given human image. This model is trained to disentangle human pose, body shape, clothing geometry, and texture. The estimated Gaussians can be animated without post-processing. We conduct comprehensive experiments to validate the effectiveness of the proposed dataset and method. Our model demonstrates the ability to efficiently reconstruct photorealistic humans at 1K resolution from a single input image using a single GPU instantly. Additionally, it seamlessly supports various applications, as well as shape and texture editing tasks.
Figure 1: This work introduces (a) IDOL, a feed-forward, single-image human reconstruction framework that is fast, high-fidelity, and generalizable; (b) Utilizing the proposed Large Generated Human Multi-View Dataset consisting of $100K$ multi-view subjects, our method exhibits exceptional generalizability in handling diverse human shapes, cross-domain data, severe viewpoints, and occlusions; (c) With a uniform structured representation, the avatars can be directly animatable and easily editable.
footnotetext: Work done during the internship at Tencent by Yiyu Zhuang, Jiaxi Lv, and Wenhao.footnotetext: Equal contributionsfootnotetext: Corresponding authors
1 Introduction
Reconstructing 3D clothed avatars from a single image is crucial in user-friendly virtual reality, gaming, and 3D content creation. This task involves mapping 2D images to 3D models, a highly ill-posed and challenging problem due to the complexity of clothing and the diversity of human poses. Learning-based methods trained on publicly available 3D human models [74, 14, 59, 21, 20] have improved reconstruction quality, but their generalization and quality remain limited [49, 50, 75, 60, 61, 73, 67]. Techniques that integrate parametric body estimation [26, 75, 60], loop optimization [60, 61, 73], and diffusion model priors [33, 57, 2] enhance performance but often result in slower and fragile reconstructions. Recently, large generic image-to-3D reconstruction and generation models [55, 17, 54, 62, 58, 32], leveraging large-scale datasets [11, 12, 40] or pre-trained diffusion models [47, 48, 3], boost capabilities but still struggle with real-life human reconstruction due to the scarcity of photorealistic 3D human data.
Back to the first principle, we rethink existing data, models, and representations. First, current data acquisition is limited by the photographing or scanning of individual subjects, making them inherently unscalable. For instance, the largest publicly available human dataset, MVHumanNet [59], contains only thousands of subjects and clothing variations, far from the dataset scale required to achieve robust model generalization across diverse input images. Second, in the realm of animatable reconstruction models, existing approaches predominantly rely on either reconstruct-and-rig methods that necessitate manual post-processing or entangle human parametric models (e.g., SMPL(-X) [39, 43]) for optimization. However, disentangling human pose and body shape, clothed geometry, and texture could simplify the learning process, avoid error accumulation in parameter estimation, and improve efficiency. Furthermore, when focusing solely on human reconstruction, the incorporation of multi-view image generation or refinement models warrants careful consideration [62, 45, 57]. While these models can introduce finer details, they may also lead to inconsistencies across different views. Third, the community of 3D vision develops various representations tailored to application-specific requirements and technological advancements [28, 41]. For human-centric applications, ideal representations should be well-structured and expressive to facilitate easy rigging and editing [35, 23], as well as capable of high-resolution and fast rendering [28] to enhance both functionality and realism.
In this work, we introduce a scalable pipeline for training a simple yet efficient feed-forward model for instant photorealistic human reconstruction. We present HuGe100K, a large-scale dataset comprising over 2.4M high-resolution ($896\times 640$) multi-view images of 100K diverse subjects. To facilitate comprehensive 3D human reconstruction, we develop a scalable data creation pipeline that integrates synthetic and real-world data, ensuring a wide range of attributes such as age, shape, clothing, race, and gender. Building upon this dataset, we introduce a novel feed-forward transformer model IDOL that efficiently predicts 3D human avatars with photorealistic textures and accurate geometry. Our approach leverages a pretrained encoder [29] and a transformer-based backbone for feature extraction and fusion, enabling instant reconstruction (within 1 second on an A100 GPU) without relying on generative models [73] for refinement. By integrating a uniform representation for the 3D human, our model achieves enhanced texture completion and ensures that the reconstructed humans are naturally animatable. Extensive evaluations demonstrate that our method excels in diverse and challenging scenarios, offering superior consistency and quality compared to existing techniques. Additionally, the efficient representation of our model enables seamless applications in animation, editing, and other downstream tasks.
Our contributions can be summarized as follows:
  • We rethink single-view 3D human reconstruction from the perspectives of data, model, and representation. We introduce a scalable pipeline for training a simple yet efficient feed-forward model, named IDOL, for instant photorealistic human reconstruction.
  • We develop a data generation pipeline and present HuGe100K, a large-scale multi-view human dataset featuring diverse attributes, high-fidelity, high-resolution appearances, and a well-aligned SMPL-X model.
  • Our comprehensive study demonstrates that leveraging large-scale generated training data significantly enhances model performance and generalizability, paving the way for further scaling up 3D human reconstruction models.
2 Related Work
2.1 3D Human Datasets
High-precision 3D human models typically rely on scanning or multi-view camera acquisition systems. Scan-based datasets, such as THuman2.0 [64], Twindom [21], and 2K2K [14], provide high-fidelity 3D geometries but are limited by a small number of subjects, simple standing poses, and non-human artifacts (e.g., 2K2K). In contrast, multi-view acquisition systems [46, 37, 66, 34, 4, 8, 27, 9, 59] facilitate the collection of larger datasets with more flexible actions. However, these datasets often face challenges such as biased indoor lighting, fixed viewpoints, and limited scale. Large-scale synthetic and real datasets for generic objects, including Objaverse and MVImgNet [11, 12, 65], address open-world reconstruction but lack specificity for human models. To overcome these limitations, the proposed HuGe100K significantly scales up dataset size, increasing subject diversity by over 100 times compared to previous datasets. It emphasizes diversity in pose, shape, viewpoint, and clothing, paving the way for large-scale model training. We compare existing datasets with ours in Tab. 1.
Type Dataset #Frames IDs SMPL(-X)
3D Scans THuman [74] - 200
THuman2.1 [64] - 2500
2K2K [14] - 2050
X-Avatar [52] 35.5K 20
Multi-view Images ZJU-MoCap [46] 180K 10
Neural Actor [37] 250K 8
HUMBI [66] 26M 772
AIST++ [34] 10.1M 30
HuMMan [4] 60M 1000
GeneBody [8] 2.95M 50
ActorsHQ [27] 40K 8
DNA-Rendering [9] 67.5M 500
MVHumanNet [59] 645.1M 4500
Ours HuGe100K $>$ 2.4M 100K
Table 1: Comparisons of related datasets. HuGe100K is a large-scale generated multi-view human dataset with $100K$ diverse high-fidelity humans.
2.2 Single-Image Human Reconstruction
For clothed 3D reconstruction methods, implicit representation methods such as PIFU [49], PIFU-HD [50], ARCH [26, 16], and PaMIR [75] have been widely adopted. To enhance reconstruction quality and generalizability, loop optimization techniques integrate implicit representations with explicit or hybrid human priors for improved robustness [60, 61, 71, 2]. Similarly, GTA [72] and SIFU [73] entangle SMPL priors and side-view conditioned features to enhance feature representation. However, these approaches rely heavily on the accuracy of SMPL estimates and are computationally expensive, typically requiring several minutes for inference. Moreover, the resulting 3D representations are not drivable. Recent advancements focus on recovering animatable 3D humans from single images [25, 24, 10]. Some methods benefit from the pre-trained diffusion models or large reconstruction models [45, 57]. Nevertheless, all methods are constrained by the limitations of training datasets, resulting in suboptimal texture detail and limited generalizability. The latest work, E3Gen [69], combines UV maps and Gaussian splatting [28] to directly generate Gaussian attribute maps in UV space from images, which does not support arbitrary image inputs.
To generate consistent human images or videos given a single image, recent diffusion video models animate a static image with pose-controllable video conditions[22, 63, 56, 6, 56, 6, 77]. Specifically, Champ [77] incorporate several rendered maps obtained from SMPL sequences, alongside skeleton-based motion guidance, to enrich the conditions to the latent diffusion model. However, they all fail to generate a 360-degree video and animate precise expressive SMPL-X-based humans due to insufficient training data.
Figure 2: Pipeline for constructing our HuGe100K. Diverse attribute combinations from GPT-4 templates create text prompts, generating synthetic images via FLUX, combined with real images from DeepFashion. SMPL-X fitting produces multi-view pose sequences with 360-degree rotations and diverse animatable motions. MVChamp then converts these sequences into multi-view images, ensuring 3D consistency in the dataset.
3 Dataset Creation
Drawing inspiration from the latest advancements in generalizable large models, our core insight is to develop a large-scale reconstruction model with exceptional generalization capabilities. The crux of this endeavor lies in creating a comprehensive and high-quality digital human dataset. In this section, we introduce the creation of a large-scale human-centric generated dataset comprising over 2.4M high-resolution (896 × 640) multi-view images from more than 100K diverse subjects.
As shown in Fig. 2, the data generation pipeline consists of two stages. Firstly, diverse text prompts are generated by large language models incorporating various human-centric attributes, and high-quality images are synthesized using text-to-image generation models (Sec. 3.1). Secondly, we train a multi-view video generation model, MVChamp, conditioned on rendered full-body motions. With this, a large-scale multi-view images dataset is established by feeding both the synthesized and captured images (Sec. 3.2). Lastly, the statistics and characteristics of the HuGe100K are demonstrated in Sec. 3.3.
3.1 Image Collection and Generation
Given the limited amount of existing datasets [59, 4, 46, 64] and legal concerns regarding portrait and copyright issues, we propose leveraging an image generation model to construct a large-scale, diverse, and high-fidelity dataset of full-body human images. Using the latest text-to-image model Flux [18], we design descriptive prompts based on the required attributes for human figures, ensuring diversity across area, clothing, body shape, age, and gender shown in Fig. 2(i). To avoid long-tail attribute distributions and ensure uniform coverage of traits, we apply uniform sampling in attribute selection for the image generation process. This approach generated over 100K images in total. However, due to issues of visual non-compliance and high similarity among some images, we conducted manual filtering and retained a final set of 90K high-quality images. Additionally, we combine 90K synthetic images and 10K real-life images from DeepFashion [38].
3.2 Multi-view Image Animation and Generation
Image-based animating models [6, 77, 22, 63, 70] struggle to achieve generation of 360-consistent and diverse human videos conditioned on pose sequences (e.g., 2D poses and 3D SMPL(-X) conditions). To address this limitation, we re-train a state-of-the-art video generative model, Champ [77], to obtain a multi-view consistent generative model, MVChamp. Firstly, We collect a curated dataset of over 100K single-person videos with various motions (e.g., dancing), including approximately 20K videos involving the action of turning around, to fine-tune the model and enhance its generalizability for diverse inputs and motions. Secondly, to further improve 3D consistency[13], we utilize scanned models from THuman 2.1 to generate uniformly distributed views via Blender. From these, we select a subset of 24 views that are evenly spaced around the full 360° rotation to fine-tune MVChamp’s temporal layers using diffusion loss. Lastly, to support high-quality whole-body animation (i.e., body and hands), we introduce precise SMPL-X estimation using NLF [51] for body shape and pose estimation. Additionally, we employ HaMeR [44] for hand estimation, rendering the corresponding hand template into depth maps as additional control signals.
After these training processes, MVChamp generates multi-view images conditioned on SMPL-X parametric model. The pose condition is set to an “A-Pose” 80% of the time and a random pose from dance videos (via SMPLer-X [5]) 20% of the time, balancing pose variability. To address discrepancies between the first and last frames, we propose a Temporal Shift Denoising Strategy: during the denoising steps, we shift latent inputs and pose signals temporally, moving the last frame to the first. This improves continuity between frames without additional inference costs. Finally, FaceFusion [19] is utilized to enhance facial details.
Figure 3: A paired example from the proposed HuGe100K Dataset. For each reference image, we generate a set of multi-view images using an estimated shape and a specific pose. The figure shows the pose is well-aligned.
3.3 Data Statistics and Characteristics
Tab. 1 compares our dataset HuGe100K to other 3D human datasets regarding scale, diversity, and consistency. With over 100K human identities and 20K poses, HuGe100K offers balanced diversity across dimensions such as area, clothing, body shape, age, and gender (see Fig. 1). Each sample includes a reference image, SMPL-X estimates, 24 uniformly sampled multi-view images, camera parameters, and SMPL-X data (see Fig. 3).
Figure 4: The architecture of IDOL, a full-differentiable transformer-based framework for reconstructing animatable 3D human from a single image. The model integrates a high-resolution ($1024\times 1024$) encoder [29] and fuses image tokens with learnable UV tokens through the UV-Alignment Transformer. A UV Decoder predicts Gaussian attribute maps as intermediate representations, capturing the human’s geometry and appearance in a structured 2D UV space defined by the SMPL-X model. These maps, in conjunction with the SMPL-X model, represent a 3D human avatar in a canonical space, which can be animated using linear blend skinning (LBS). The model is optimized using multi-view images with diverse poses and identities, learning to disentangle pose, appearance, and shape.
4 Large Human Reconstruction Model
In this section, we present the large-scale human reconstruction model, named IDOL. The overview pipeline is shown in Fig. 4. In Sec. 4.1, we describe the animatable human representation. Sec. 4.2 details the network architecture of IDOL, while Sec. 4.3 explains how we trained the model in an end-to-end manner using the multi-view image dataset.
4.1 Animatable Human Representation
Similarly to previous work [69, 23, 31], IDOL leverages 3D Gaussian Splatting [28] in conjunction with SMPL-X for 3D human representation, aiming to address the challenges of real-time rendering and accurate animation of human avatars. Specifically, each Gaussian primitive $\mathcal{G}_{k}$ is characterized by: $\mathcal{G}_{k}=\left\{\mu_{k},\alpha_{k},\mathbf{r}_{k},\mathbf{s}_{k},\mathbf{c}_{k}\right\},$ where $\mu_{k}$ is the 3D position of the Gaussian, $\alpha_{k}$ is opacity, $\mathbf{r}_{k}$ is the rotation, $\mathbf{s}_{k}$ is the scale, and $\mathbf{c}_{k}$ is the color.
3D Gaussian Human.
Directly predicting all 3D Gaussian primitives is computationally intensive. Instead, IDOL leverages the predefined 2D UV space of the SMPL-X model to transform the 3D representation task into a more manageable 2D problem. Initially, IDOL predicts Gaussian attribute maps that encode the offset values {{\${\delta_{\mathbf{\mu}_{k}},\delta_{\mathbf{r}k},\delta{\mathbf{s}_{k}}}$}}\}}, as well as the color $\mathbf{c}_{k}$ and opacity $\alpha_{k}$ for each Gaussian primitive $\mathcal{G}_{k}$. Similarly to $E^{3}Gen$ [69], the position $\mathbf{\mu}_{k}$, scale $\mathbf{s}_{k}$, and rotation $\mathbf{r}_{k}$ of each Gaussian primitive $\mathcal{G}_{k}$ is modeled relative to its SMPL-X vertex as follows: $\mathbf{\mu}_{k}=\hat{\mathbf{\mu}}_{k}+\delta_{\mathbf{\mu}_{k}},\mathbf{s}_{k}=\hat{\mathbf{s}}_{k}\cdot\delta_{\mathbf{s}_{k}},\mathbf{r}_{k}=\hat{\mathbf{r}}_{k}\cdot\delta_{\mathbf{r}_{k}},$ where $\hat{\mathbf{\mu}}_{k}$, $\hat{\mathbf{s}}_{k}$, and $\hat{\mathbf{r}}_{k}$ are the initial values based on the SMPL-X model. The color $\mathbf{c}_{k}$ is assigned using RGB values, and the opacity $\alpha_{k}$ is set to $1$, indicating full opacity. This unified UV space significantly reduces computational complexity and fully exploits the geometric and semantic priors provided by the SMPL-X model. By modeling Gaussian primitives relative to SMPL-X vertices, IDOL ensures semantic consistency across corresponding body parts of diverse avatars, thereby enhancing the model’s generalization capability across various reference images.
Animation and Rendering.
Given a target pose, we calculate the transformation of each human joint using predefined kinematic relationships. The transformation of each Gaussian primitive is performed using a forward skinning technique based on LBS. Specifically, the position of each Gaussian primitive $\mathcal{G}_{k}$ is transformed as follows: $\mathbf{\mu}_{k}^{\prime}=\sum_{i=1}^{n_{b}}w_{i}\mathbf{B}_{i}\mathbf{\mu}_{k}$. Additionally, the rotation matrix $\mathbf{R}_{k}$ is updated by: $\mathbf{R}_{k}^{\prime}=\mathbf{T}_{k}^{1:3,1:3}\mathbf{R}_{k},$ where $\mathbf{T}_{k}=\sum_{i=1}^{n_{b}}w_{i}\mathbf{B}_{i},$ and $\mathbf{R}_{k}$ is the rotation matrix representation of the rotation angle $\mathbf{r}_{k}$. Here, $n_{b}$ is the number of joints, $\mathbf{B}_{i}$ is the transformation matrix for each joint, and $w_{i}$ represents the skinning weights, indicating the influence of each joint’s motion on the Gaussian primitive’s position $\mathbf{\mu}_{k}$.
To estimate the skinning weights for each $\mathcal{G}_{k}$, we compute the skinning weight field for body parts using a pre-defined low-resolution volume[7]. For the smaller regions, such as the hands and face, which exhibit minimal topology changes, their skinning weights are calculated by interpolating from the SMPL-X template using barycentric coordinates. This approach supports modeling large-scale topology changes among different identities, including substantial changes introduced by clothing. In addition, it helps stabilize the convergence for areas like the fingers and face, which experience fewer topological changes compared to the SMPL-X template.
4.2 Network Structure
As illustrated in Fig. 4, IDOL is a full-differentiable transformer-based framework for reconstructing animatable 3D human from a single image.
High-resolution Image Encoder.
Higher resolution of the input image results in a high-quality reconstruction. However, previous works suffer from low-resolution ViT-based encoders, such as DINOv2 [42], which support the $448\times 448$ resolution. To fully leverage the resolution of the HuGe100K dataset, we further adopt a high-resolution human foundation model, Sapiens [29], to encode the $1024\times 1024$ resolution images into patch-wise feature tokens, formulated as: $\mathbf{F}=\{\mathbf{f}_{i}\}^{n}_{i=1}\in\mathbb{R}^{d_{E}},$ where $i$ denotes the $i$-th image patch, and $d_{E}$ is the channel length of each token. The Sapiens model is pretrained on 300 million in-the-wild human images using the Masked Autoencoder (MAE) framework [15], enabling it to excel in preserving fine-grained details and capturing diverse human poses and appearances, making it highly effective for high-resolution human image feature extraction.
UV-Alignment Transformer.
To map irregular and diverse reference images onto regular UV feature maps, a UV-Alignment Transformer is employed to align learnable spatial-positional UV tokens $\mathbf{Q}^{0}$ with reference image features $\mathbf{F}$. Specifically, the UV-Alignment Transformer concatenates $\mathbf{F}$ from Spaiens with $\mathbf{Q}^{0}$, and aggregates and refines features through $D$ transformer blocks, producing an enhanced representation $\mathbf{Q}^{D}$. Each transformer block utilizes self-attention mechanisms, enabling the model to capture complex relationships among the input tokens and impute missing parts using correlated tokens.
UV Decoder.
The UV tokens $\mathbf{Q}^{D}$ are reshaped and decoded into Gaussian attribute maps, capturing the human’s geometry and appearance within the structured 2D UV space as shown in Sec. 4.1. To preserve and enhance fine details in the decoded Gaussian attribute maps, a Convolutional Neural Network is employed for spatial up-sampling.
4.3 Training Objectives
IDOL reconstructs 3D human by predicting Gaussian attribute UV maps in a single forward pass, offering significant advantages in inference speed and enabling end-to-end training with multi-view images. For each sample, we select a front view as the reference image $\mathbf{I}_{\text{ref}}$, along with a random set of generated multi-view images $\{{\mathbf{I}_{\text{gt},i}}\}_{i=1}^{N}$ and their corresponding SMPL-X parameters and camera for supervision. It takes $\mathbf{I}_{\text{ref}}$ as input to generate the 3D human, using differentiable rendering to produce the multi-view images $\{{\mathbf{I}_{\text{pred},i}}\}_{i=1}^{N}$. The loss function is defined as follows:
$\mathcal{L}=\sum_{i=1}^{N}\left(\left\lVert\mathbf{I}_{\text{gt},i}-\mathbf{I}_{\text{pred},i}\right\rVert^{2}+\lambda L_{vgg}(\mathbf{I}_{\text{gt},i},\mathbf{I}_{\text{pred},i})\right),$ (1)
where $\lambda$ controls the balance between the mean square error loss and the perceptual loss $L_{vgg}$.
5 Experiments
Figure 5: Qualitative results of our MVChamp ablation study (left) and comparison experiment (right).
5.1 Experiments on Dataset Creation
Fig. 5 (left) presents the qualitative results of the ablation study on our data generation method. Specifically, Fig. 5 (a) and Fig. 5 (b) demonstrate how hand refinement and face swapping effectively enhance the quality of the generated hands and faces. Fig. 5 (c) highlights the importance of fine-tuning the 3D dataset Thuman 2.1 for improving the 3D consistency of MVChamp, while Fig. 5 (d) illustrates that the Temporal Shift Denoising strategy, which involves cycling the first and last frames, improves the consistency between the first and last frames.
Fig. 5 (right) compares our MVChamp’s generation quality against previous human image animation models (e.g., Champ, MimicMotion [70]) and general multi-view video generation model (e.g., SV3D [55]). The issues of these methods are highlighted in the figure. Champ struggles to generate plausible hands and heads in multi-view scenarios. MimicMotion not only fails to produce realistic shoes but also has difficulty preserving the identity, especially for the anime image. SV3D produces low-quality multi-view human generation due to the lack of 3D prior knowledge of the human body. In contrast, our model generates consistent, high-quality results across views.
5.2 Comparison of Reconstruction Model
Implementation Details.
For the encoder, we employed the pre-trained Sapiens-1B model[29] and kept its weights frozen during training. We then defined a transformer-based framework with a parameter size of $0.5$ billion to perform feature fusion. Additionally, we densify the SMPL-X vertex set by sampling approximately 200,000 vertices to map attributes and represent the 3D human model effectively.
Dataset and Metrics.
We train IDOL on a dataset consisting of generated multi-view images from HuGe100K and rendered images from THuman 2.1 [74]. For THuman 2.1, we render 72 view images at a resolution of $896\times 640$ for each scan, with rendering views uniformly distributed. To evaluate performance quantitatively, we reserve the last 50 cases from both HuGe100K and THuman 2.1 as the testing set. We use the most frontal image as the reference image and the remaining images as ground-truth data for evaluation. The ground-truth camera parameters and SMPL-X parameters are provided for all methods. The renderings of the reconstructed model are then compared to their corresponding ground truth images with several metrics, including Mean Squared Error (MSE), Learned Perceptual Image Patch Similarity (LPIPS) [68], Peak Signal-to-Noise Ratio (PSNR). More details on implementation can be found in the supplementary material.
Figure 6: Comparisons on (a) the upper: novel-view synthesis given a single image, and (b) the lower: our animated results.
Baselines.
We compare IDOL with three baseline categories in the Single-Image Human Reconstruction task. The first category comprises methods based on loop optimization or pixel-alignment modules, including GTA [72] and SIFU [73]. These methods focus on refining human reconstruction through iterative optimization or predicting 3D geometry from pixel-aligned features.
The second category encompasses large-scale generic reconstruction networks, represented by LGM [54]. These models are known for their ability to handle large datasets and their scalability, offering advantages in terms of fast reference speeds and large output resolutions.
The third category involves optimization-based 3D generation methods using score distillation sampling (SDS), exemplified by DreamGaussian [53]. These methods leverage priors from 2D diffusion models to distill 3D objects. DreamGaussian accelerates convergence by progressively densifying 3D Gaussians, significantly reducing the reconstruction time. However, it still takes approximately two minutes to reconstruct a single object.
Quantitative comparison.
As shown in Tab 2, our method outperforms all baselines in all metrics. We attribute this superior performance to the large-scale dataset HuGe100K and the design of our large-scale reconstruction model IDOL, which allows for more effective training and improved synthesis of appearance results. Despite this, we note that SIFU and GTA report a lower metric than what we expected in our test settings. While we provide accurate ground-truth SMPL-X parameters and camera settings, the ideal orthographic projection required by the pixel-alignment modules in SIFU and GTA is not well-suited to our perspective projection model, where the camera focus ranges from $35$ to $80$. This mismatch leads to a misalignment between the rendered images and the ground truth, adversely affecting their performance metrics. In fact, many real-life photographs are taken with medium to short focal lengths and cannot be approximated using orthographic projection, which is a less noticeable drawback of SIFU and GTA. Additionally, SIFU and GTA, trained with images at a resolution of $512\times 512$, struggle to synthesize detailed textures, particularly in the invisible areas of the reference images. This limitation is primarily due to the lack of comprehensiveness and diversity in their training datasets, which restricts their performance in generating invisible aspects of the images.
Qualitative comparison.
We perform a qualitative evaluation on an in-the-wild dataset with methods from [73, 72, 53, 54], with the results presented in Fig. 6. Our evaluation includes a subject in a complex outfit featuring a baseball cap and textured clothing, demonstrating our method’s ability to synthesize intricate textures and handle loose outfits across different views. Additional tests include out-of-domain cartoon data and large-angle side views and assessing model adaptability and viewpoint handling. IDOL consistently outperforms the baselines, which struggle with detail reproduction and texture consistency under these various conditions.
5.3 Ablation Study on Reconstruction Model
We assess the impact of proposed components by removing them individually from IDOL. As shown in Fig. 7, replacing the Sapiens [29] encoder with DINO v2 [42] (w/o Sapiens) reduces detail quality, resulting in less realistic textures and folds. Excluding the HuGe100K dataset (w/o HuGe100K) causes significant distortion, including blurred details and color bleeding. The complete model, with all components, produces the most realistic and detailed results. These findings highlight the critical roles of both the Sapiens encoder and the HuGe100K dataset in achieving high-quality avatar generation.
Figure 7: Qualitative Results of Ablation Study of IDOL.
Method MSE $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
SIFU [73] 0.042 14.204 1.612
GTA [72] 0.041 14.282 1.629
Ours-w/o HuGe100K 0.017 19.225 1.326
Ours-full 0.008 21.673 1.138
Table 2: Evaluation of Comparison and Ablation Experiments.
5.4 Downstream Applications
Figure 8: Controllable Avatar Editing: (a) texture editing; (b) body shape editing.
IDOL reconstructs an avatar by combining Gaussian attribute maps with the SMPL-X model, enabling users to modify the avatar’s appearance by editing UV texture maps and to control body shape by adjusting SMPL-X shape parameters. Fig. 8(a) illustrates the effect of texture editing on clothing patterns, while Fig. 8(b) demonstrates shape editing to adjust the avatar’s body size. This approach provides a high degree of controllability over both the avatar’s appearance and body shape.
6 Conclusions and Limitations
In conclusion, our work has made significant strides in creating an animatable 3D human from a single image. We introduced a scalable pipeline for training a simple yet efficient feed-forward model, incorporating a dataset generation framework, a large-scale dataset HuGe100K, and a scalable reconstruction transformer model, IDOL. This model efficiently reconstructs photorealistic 3D humans in under a second and demonstrates versatility across various applications. This model can efficiently reconstruct photorealistic humans in less than a second and is versatile enough to support various applications.
However, there are limitations. Due to the constraints of the video model used, we can only synthesize single-frame images from fixed viewpoints. Future work could consider generating longer motion sequences. Additionally, our dataset lacks the generation of data involving multiple people interacting or people interacting with objects. Our method also struggles with handling half-body inputs. Future efforts could focus on enhancing data generation strategies to improve model performance.
Supplementary Material
In this supplementary material, we provide additional details and visualizations to support the claims made in our main paper. Sec. A provides further details on the HuGe100K dataset, including visualizations, important statistics, and the methodology to enhance the 3D consistency and diversity of multi-view images. Sec. B describes the training procedure and setup for our proposed method, IDOL. Sec. C presents additional experimental results, including comparison tables and results from the user study.
Figure 9: The visualization of the reenactment.
Appendix A More Details of HuGe100K Data
This section provides a more detailed explanation of the HuGe100K dataset generation process, along with additional visualizations. Sec A.1 describes our approach to improving the 3D consistency of MVChamp during training, while Sec A.2 presents the prompt template and attribute set used to generate reference images, as well as the generation process with MVChamp. Sec. A.3 demonstrates more visualization of the dataset.
A.1 Improving 3D Consistency of Image Animation
Champ [77] is one of the state-of-the-art Human Image Animation models, enhanced by multiple conditions rendered from DWPose and SMPL. We employ a two-stage training process to enhance the 3D consistency of the Champ model for human multi-view synthetic, referred to as MVChamp.
Fine-tuning Champ on Large-scale Human Videos with Whole-body Conditions
To enable Champ to learn more human 3D prior knowledge, we curate a dataset of approximately 100K dance videos for fine-tuning, of which around 20K explicitly contain human turning motions. Full parameter training of MVChamp on such a large dance dataset effectively enhances its understanding of human 3D prior knowledge. Additionally, we employ HaMeR [44], a state-of-the-art model for 3D hand reconstruction, to specifically reconstruct hand poses from images. These reconstructed hand poses are rendered into depth maps and used as an additional pose control signal for precise whole-body reconstruction and animation.
Fine-Tuning Temporal Blocks on 3D Human Dataset
We use the open-source scanned dataset THuman 2.1 [74], rendered in Blender, to produce 24 uniformly sampled views along the horizontal dimension to fine-tune the temporal layers of MVChamp using standard diffusion loss.
Improving Temporal Consistency from the First to Last Frames
Although the MVChamp model generates highly continuous multi-view images between adjacent frames, significant discrepancies remain between the first and last views, even though these two views are continuous in content. This issue likely arises from the model’s emphasis during training on ensuring continuity between adjacent frames while neglecting the larger temporal gap between the first and last frames. Thus, we propose the Temporal Shift Denoising Strategy to address this issue. During each denoising step, we shift the current latent inputs and pose condition signals along the temporal axis, moving the latent inputs and pose condition of the last frame to the first frame. This strategy ensures that each frame can access contextual information during most of the denoising steps, effectively eliminating discrepancies between the first and last frames at the same inference cost.
A.2 Generating Balanced and Diverse Images
Balanced, diverse, high-quality, high-resolution, and full-body images are scarce in existing human-centric datasets, and they are challenging to collect on the Internet due to copyright and portrait rights issues. Therefore, we mix the real-life images and generate photorealistic images to obtain the large-scale quantity and high-quality images. Specifically, we extract approximately 10,000 real-life images from the open-source dataset DeepFashion [38] and use Flux [18], a state-of-the-art text-to-image model, to generate balanced and diverse human reference images. We ensure balance and diversity across five dimensions during image generation: area, clothing, body shape, age and gender. Each dimension value is randomly selected from a large set of options generated by GPT-4 [1], with prompt templates as follows: Front view, full-body pose of a {age} old {body shape} {area} {gender} wearing {clothing} and visible hands. He/She stands against a white background, evenly lit. Ultimately, we collect a total of 100,000 balanced and diverse full-body human reference images.
For each dimension, the possible options are as follows:
  1. 1.
    Area: United States, Canada, Mexico, Guatemala, Cuba, Brazil, Argentina, Colombia, Chile, Peru, United Kingdom, Germany, France, Italy, Spain, Netherlands, Belgium, Switzerland, Poland, Sweden, Nigeria, Egypt, South Africa, Kenya, Morocco, Ghana, Tanzania, Ethiopia, Uganda, Algeria, Saudi Arabia, Iran, Turkey, Israel, United Arab Emirates, Qatar, Kuwait, Jordan, Oman, Lebanon, Kazakhstan, Uzbekistan, Turkmenistan, Kyrgyzstan, Tajikistan, India, Pakistan, Bangladesh, Sri Lanka, Nepal, Bhutan, China, Japan, South Korea, Mongolia, North Korea, Indonesia, Thailand, Vietnam, Malaysia, Philippines, Singapore, Myanmar, Cambodia, Laos, Brunei, Australia, New Zealand, Papua New Guinea, Fiji, Solomon Islands, Jamaica, Haiti, Dominican Republic, Puerto Rico, Trinidad and Tobago, Panama, Costa Rica, Nicaragua, Honduras, El Salvador, Belize, etc.
  2. 2.
    Clothing: T-shirts, Jeans, Casual pants, Dresses, Shorts, Tank tops, Sweaters, Cardigans, Jumpsuits, Hoodies, Suits, Business shirts, Formal skirts, Dress pants, Blazers, Tie, Waistcoats, Formal shoes, Briefcases, Leather belts, Sport shirts, Fitness clothes, Sports shoes, Tracksuits, Gym shorts, Leggings, Swimwear, Cycling gear, Compression wear, Evening gowns, Tuxedos, Long dresses, Tailcoats, Cocktail dresses, Party wear, Ceremonial suits, Ball gowns, Dress shoes, Fine jewelry, Hiking clothes, Waterproof jackets, Thermal wear, Camping gear, Fishing vests, Hunting apparel, Snowboarding pants, Rain boots, Cotton shirts, Linen dresses, Chiffon blouses, Sandals, Sunglasses, Short sleeves, Beachwear, Crop tops, Wool coats, Thick cotton sweaters, Fur jackets, Beanies, Boots, Gloves, Scarves, Thermal leggings, Padded parkas, Insulated boots, Hanfu, Kimono, Sari, African tribal dresses, Scottish kilts, Bavarian lederhosen, Moroccan kaftans, Hawaiian shirts, Russian ushankas, Streetwear, Avant-garde designs, Fusion wear, Boho chic, Minimalist styles, High fashion, Urban outfits, Eco-friendly clothing, Techwear, Nurse uniforms, Firefighter gear, Construction vests, Police uniforms, Military boots, Lab coats, Coveralls, Military uniforms, Academic gowns, Judicial robes, Clerical vestments, Diplomatic suits, Regalia, etc.
  3. 3.
    Body shape: Slight, Lean, Petite, Athletic, Fit, Average, Built, Buff, Bodybuilder, Full-figured, Stocky, Large.
  4. 4.
    Age: 20–30 years, 30–40 years, 40–50 years, 50–60 years, 60–70 years, 70–80 years, 80–90 years.
  5. 5.
    Gender: Female and male.
A.3 Additional Visualization
Fig. 11 shows the diversity of reference images generated using our prompt template and attribute set. Fig.12 and Fig. 13 illustrate the multi-view images under diverse poses generated by our MVChamp.
A.4 Application: Human Video Reenactment
The goal of this application is to replace a person in a reference video with a new identity while preserving the background and pose. Given a reference image that provides the target identity, and a reference video that provides the pose and background of the original person, the task is to seamlessly swap the person in the video while maintaining the integrity of the scene. We visualize the results in Fig. 9.
To achieve this, we follow a multi-step process:
Identity Reconstruction: The IDOL model is used to reconstruct an animatable 3D human from the reference image. This model generates a highly detailed and realistic representation of the target identity, allowing us to manipulate the avatar to match various poses.
Background Inpainting: The video inpainting process restores the regions of the video frame where the original person has been replaced, ensuring a seamless background. It involves detecting and tracking the target area using a segmentation method, which is initialized and refined by the widely used zero-shot segmentation model, Segment Anything Model (SAM)[30]. Once the target area is segmented and tracked, the remaining regions are completed using the video inpainting method, ProPainter[76], ensuring the background is seamlessly restored with no traces of the replaced identity.
Pose Animation: The target pose is extracted from the reference video [36, 5], and the reconstructed human model is animated to match this pose. The IDOL model provides precise control over the 3D human’s pose, including fine details such as finger movements, allowing it to adapt dynamically to the reference video’s actions. After animating the 3D human, we render it into the target view and seamlessly blend it with the background.
Utilizing IDOL, our process offers an efficient and high-quality solution for identity replacement in videos, providing greater stability and lower computational cost compared to 2D-based approaches [22, 77]. This opens up new possibilities for digital content creation and interactive media applications.
A.5 Representation Comparisons
To further illustrate the differences between our method and previous approaches, we provide a comparison in Fig. 10. Below, we explain the key differences:
Comparison to PIFU: PIFu predicts the 3D human shape directly from a given image without leveraging a parametric model prior. While effective for simple cases, it often lacks robustness and precision, particularly when handling challenging poses or incomplete observations [60].
Comparison to GTA/SIFU: GTA and SIFU utilize loop optimization [60, 61] to align the reconstructed output with SMPL models. While this alignment step is crucial for pixel-aligned operations [49], it introduces several significant drawbacks:
- High computational cost: Loop optimization requires multiple iterations, adding several minutes of processing time. Additionally, it depends on the estimation of intermediate representations such as masks, normals, and skeletons.
- Error accumulation: Misalignments during optimization can accumulate over iterations, degrading the quality of the final 3D human reconstruction.
Our Approach: In contrast, our method adopts a direct and efficient pipeline: We extract image features using a large-scale encoder [29], which captures rich and detailed visual information. We then predict the 3D human shape and appearance in a uniform space, directly providing the 3D human reconstruction along with the estimated SMPL-X parameters.
By decoupling feature extraction from SMPL-X-based 3D prediction, our approach avoids the error accumulation inherent in optimization-based methods. When pose information is unnecessary, our method relies primarily on body shape estimation, reducing the dependency on precise pose alignment. Furthermore, our method supports direct animation and editing (e.g., shape and texture), unlocking additional applications and expanding its potential value in digital content creation.
Figure 10: Visualization of different approaches for 3D human reconstruction. Unlike PIFu, which directly predicts the 3D human without a parametric prior, and GTA/SIFU, which relies on computationally expensive loop optimization for SMPL alignment, our IDOL method leverages SMPL-X as a prior. This enables more robust and accurate reconstruction while avoiding the pitfalls of error accumulation. Furthermore, our method supports direct animation and editing, enabling additional applications in digital content creation.
Figure 11: Visualization of diverse images generated by Flux [18].
Figure 12: Visualization of examples from HuGe100K, where the images are generated by Flux and used to generate multi-view images.
Figure 13: Visualization of examples from HuGe100K, where the images are derived from the DeepFashion [38] dataset and used to generate multi-view images.
Appendix B More Details of IDOL
In this section, we describe the training setup and methodology for our proposed method, IDOL.
B.1 Implement Details
Our models are trained on a cluster of $32$ NVIDIA H100 GPUs for approximately 1 day, with a batch size of $32$. The optimization is performed using the Adam optimizer with a learning rate of $45e-$ 4. A warm-up schedule of $3,000$ steps is employed to stabilize training in the initial stages.
The training loss function is a weighted combination of VGG perceptual loss and Mean Squared Error, balanced with a $1:1$ ratio. This loss formulation ensures both perceptual quality and pixel-wise accuracy.
B.2 Network Architecture
The proposed network consists of a multi-stage structure designed for high-dimensional feature extraction and reconstruction tasks. The primary components include the pre-trained encoder, UV-Alignment Transformer, and UV decoder. For the encoder, we utilize the large-scale model Sapiens [29] to extract and tokenize human features from the input image.
UV-Alignment Transformer. The neck module employs a hierarchical design inspired by recent advancements in vision transformer architectures [29], featuring a decoder embedding layer with a width of $1536$ and $16$ transformer layers. Each transformer encoder layer consists of the following components:
  1. 1.
    A layer normalization operation for input stabilization, enhancing training dynamics, and preventing gradient instability.
  2. 2.
    A multi-head self-attention mechanism that maps inputs into query, key, and value representations, followed by a linear projection layer to integrate attention outputs. This process is regularized through dropout for improved generalization and further normalized to ensure consistent feature scales.
  3. 3.
    A feed-forward network (FFN) composed of two dense layers with a GeLU activation function applied between them. The FFN architecture is complemented by intermediate normalization layers to enhance stability and improve optimization convergence.
UV Decoder. The decoder begins by reshaping tokens into a 2D feature map of $64\times 64$ resolution. It employs a hierarchical upsampling and convolutional strategy to progressively refine and synthesize outputs. The upsampling mechanism uses transposed convolutional layers to increase spatial resolution, with each stage incorporating normalization and non-linear activation for stable feature transformations. Specifically:
  1. 1.
    Upsampling Blocks: The decoder incorporates multiple transposed convolutional layers, which double the spatial resolution at each stage. Instance normalization and SiLU activations provide stable scaling and enable non-linear feature transformations.
  2. 2.
    Convolution Block: Three convolutional layers with output channels $\{128,128,32\}$ further process the features, applying instance normalization and activation functions to improve feature quality and representation.
Head Module. Following [69], we construct two distinct convolutional networks for decoding geometry and color separately. These networks progressively process feature channels, transitioning from an initial channel size of $32$ to the target parameters $\delta_{\mu_{k}},\delta_{\mathbf{s}_{k}},\delta_{\mathbf{r}_{k}}$ for geometry and $\mathbf{c}_{k}$ for color.
Appendix C Experiment
In this section, we present additional experimental results, including comparison tables and the user study. We show additional visual comparisons in Fig. 14 and Fig. 15. We compare with the reported results by Weng et al. [57] and AlBahar et al. [2].
Figure 14: More visualization for comparison in the in-the-wild cases. We compare with the reported results by HumanLRM[57].
Figure 15: More visualization for comparison in the in-the-wild cases. We compare with the reported results by HumanSGD[2].
User Study.
We conducted a user study with 20 participants via evaluating 50 cases. Participants ranked results based on face, clothing, back-view consistency, and the overall quality. The aggregated results are presented in Tab. 3, showing the superiority of our method.
Method Face Clothing Back Overall
GTA 0% 2.27% 0% 0%
SIFU 2.27% 4.55% 4.55% 4.55%
HumanLRM 45.45% 43.18% 36.36% 45.45%
Ours 52.28% 50.0% 59.09% 50%
Table 3: The user study. We evaluated IDOL on selected cases reported by HumanLRM[57], designed to highlight their strengths. Despite the selection for HumanLRM, our method achieves slightly superior performance, demonstrating greater robustness and effectiveness under comparable conditions.